OpenStudy (iwanttogotostanford):

AM I right here?: Describe how to transform the graph of f into the graph of g. f(x) = x^4 and g(x) = -x^4 - I think the answer is Reflect the graph of f across the y-axis.

OpenStudy (iwanttogotostanford):

@calculusxy @phi @YoungStudier @retirEEd

satellite73 (satellite73):

\((1,1)\) is on the first graph the second has \((1,-1)\) \((216)\) on the first , \((2,-16)\) on the second

satellite73 (satellite73):

oops i meant \((2,16)\) on the first, \((2,-16)\) on the second

OpenStudy (mathmale):

Try graphing the simpler function y=x^2. Then, graph h=-x^2. How does the graph of y become the graph of h?

OpenStudy (iwanttogotostanford):

but was I right in thinking that it flips over y axis? Im guessing no..?

satellite73 (satellite73):

no

OpenStudy (mathmate):

If you follow @mathmale 's advice, you will find the answer very quickly.

OpenStudy (iwanttogotostanford):

oh ok I did it and i figured i have to flip across the X AXIS not y

satellite73 (satellite73):

in fact \(y=x^4\) is symmetric with respect to the y axis, so if you flip it about the y axis nothing changes

satellite73 (satellite73):

yes

OpenStudy (iwanttogotostanford):

thank you! That was easy.

OpenStudy (mathmale):

So, if you rotate the graph of x^2 about the x-axis, you obtain the graph of -x^2. Now, how would you answer the problem you initially posted?

satellite73 (satellite73):

what was up is now down since it has a minus sign in front that is always true of \(y=-f(x)\) compared to \(y=f(x)\)